OpenStudy (anonymous):

y varies jointly with x and the square of n. when y = 25, x =3 and n=4. Find the value of y when x - 5 and n = 10. Round to the nearest whole number.

OpenStudy (anonymous):

260?

OpenStudy (anonymous):

oh the choices are 360 320 290 260

OpenStudy (saifoo.khan):

\[y = kx^n\]

OpenStudy (anonymous):

y=k(x^n)(y^m)

OpenStudy (saifoo.khan):

how can do write that?? y^m ??

OpenStudy (anonymous):

thats the formula they gave me :P

OpenStudy (saifoo.khan):

Your school is prolly wrong man!

OpenStudy (anonymous):

i dont know what to say...

OpenStudy (saifoo.khan):

or might be you are mixing something

OpenStudy (anonymous):

ill send you the link they gave me...

OpenStudy (saifoo.khan):

Alright.

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

25=k48 k=25/48 now y=260.47

OpenStudy (anonymous):

thank you!